OpenStudy (anonymous):

Solve the absolute value equation|4x-1|=1 answer are. X=2or x=0, x=0or x=-1/2, x=0or x=1/2, x=2or x=-1/2

OpenStudy (anonymous):

x=0 or x=1/2

OpenStudy (jdoe0001):

\(\large |4x-1|=1\implies \begin{cases} +(4x-1)=1\\ \quad \\ -(4x-1)=1 \end{cases}\)